# Revealing the building blocks of tree balance: fundamental units of the Sackin and Colless Indices

### v1

@@ Abstract
Over the past decades, more than 25 phylogenetic tree balance indices and several families of such indices have been proposed in the literature – some of which even contain infinitely many members. It is well established that different indices have different strengths and perform unequally across application scenarios. For example, power analyses have shown that the ability of an index to detect the generative model of a given phylogenetic tree varies significantly between indices. This variation in performance motivates the ongoing search for new and possibly “better” (im)balance indices. An easy way to generate a new index is to construct a compound index, e.g., a linear combination of established indices.
Two of the most prominent and widely used imbalance indices in this context are the Sackin index and the Colless index. In this study, we show that these classic indices are themselves compound in nature: they can be decomposed into more elementary components that independently satisfy the defining properties of a tree (im)balance index. We further show that the difference Colless minus Sackin results in another imbalance index that is minimized (amongst others) by all Colless minimal trees. Conversely, the difference Sackin minus Colless forms a balance index.
Finally, we compare the building blocks of which the Sackin and the Colless indices consist to these indices as well as to the stairs2 index, which is another index from the literature. Our results suggest that the elementary building blocks we identify are not only foundational to established indices but also valuable tools for analyzing disagreement among indices when comparing the balance of different trees.
@@ 1 Introduction
This variability has led to the continuing development of new indices, often with the aim of capturing specific aspects of tree balance more effectively. A common approach in the design of new indices is to combine established ones into compound indices, for instance via simple linear combinations [11, Chapter 25], [16, 4, 20]. Among the most commonly used base indices for such compound construction are the Sackin index [22, 10], which measures the total of all subtree sizes, and the Colless index [7, 8], which measures subtree size imbalances across all internal nodes. Both have been used extensively in theoretical studies and practical applications [2, 3, 25, 11, 15], and they often serve as components in the construction of new indices [11, 4, 16, 20].
@@
Building on this decomposition of the Sackin and Colless indices, we even identify two more indices, which are new to the literature and which turn out to be closely related to $N_{a}$ and $N_{b}$: the difference Colless-Sackin, $\Delta_{CS}$, forms an imbalance index that remains minimal (amongst others) for the same class of balanced trees that minimize Colless, while the difference Sackin-Colless, $\Delta_{SC}$, constitutes a balance index. Concerning all newly discovered indices, i.e., $N_{a}$, $N_{b}$, $\Delta_{CS}$ and $\Delta_{SC}$, we are able to fully characterize their extremal trees for all $n$, and in this characterization, the so-called echelon tree plays a crucial role. The echelon tree has so far mainly been known from the context of another tree balance index from the literature, namely the so-called stairs2 balance index [9]. In our study, we also present some new insight into this tree and present, for instance, the first non-recursive way to generate it.
@@ 2.1 Definitions and notation
The most important concept for the present manuscript is that of a rooted binary tree (or simply tree for short), which is a directed graph $T=(V(T),E(T))$, with vertex set $V(T)$ and edge set $E(T)$, containing precisely one vertex of in-degree zero, namely the root (denoted by $\rho$), such that for every $v\in V(T)$ there exists a unique path from $\rho$ to $v$, there are no vertices with out-degree one and such that each vertex either has has out-degree 2 (such vertices will be referred to as inner vertices) or 0 (such vertices will be referred to as leaves). For every $n\in\mathbb{N}_{\geq 1}$, we denote by $\mathcal{BT}^{*}_{n}$ the set of (isomorphism classes of) rooted binary trees with $n$ leaves.
@@
Recall that whenever you have a rooted tree $T$, you can define the depth $\delta_{v}$ of a vertex $v$ of $T$ as the length of the unique path from the root $\rho$ of $T$ to $v$. The height $h(T)$ of $T$ is then simply the maximum of all leaf depths, i.e., $h(T)=\max_{x\in V^{1}(T)}\delta_{x}$. Given a subset $X^{\prime}$ of leaves of $T$, the lowest common ancestor or lca of $X^{\prime}$ is the unique ancestor of all leaves in $X^{\prime}$ with the highest depth.
@@
We continue with the greedy from the bottom tree, denoted by $T^{gfb}_{n}$. It is the rooted binary tree with $n$ leaves that results from greedily clustering trees of minimal leaf numbers, starting with $n$ single vertices and proceeding until only one tree is left as described by [8, Algorithm 2].
@@
Another tree that will play a crucial role in our manuscript is the binary echelon tree or echelon tree for short, which is denoted by $T_{n}^{be}$. This tree was recursively defined in [9] as follows: If $n=0$, $T_{n}^{be}$ is the empty tree $T^{\emptyset}$, and if $n=1$, $T_{n}^{be}$ consists of a single vertex. Now, for $n>1$, let $k_{n}=\left\lceil\log_{2}n\right\rceil$. Then, $T_{n}^{be}$ contains a maximal pending subtree $T_{a}$ which equals $T_{k_{n}-1}^{fb}$ and a maximal pending subtree $T_{b}$ which equals $T_{n-2^{k_{n}-1}}^{be}$. Thus, the binary echelon tree is uniquely defined (up to isomorphism) for all $n$.
@@
In order to define (im)balance indices, however, we first need to consider tree shape statistics: A tree shape statistic is a function $t:\mathcal{BT}^{*}_{n}\rightarrow\mathbb{R}$ that depends only on the shape of $T$ but not on the labeling of vertices or the length of edges. Based on this, a tree imbalance index is defined as follows: A tree shape statistic $t$ is called an imbalance index if and only if
@@
Similarly, a tree shape statistic $t$ is called an balance index if and only if
@@
Following, for instance, [11], we call a balance or imbalance index local if it fulfills the following criterion: If two trees differ only in one rooted subtree, the difference between their indices is equal to the difference between the indices of these differing subtrees [11, 21]. More precisely, a balance or imbalance index $t$ is local, if $t(T)-t(T^{\prime})=t(T_{v})-t(T_{v}^{\prime})$ for $T$, $T^{\prime}\in\mathcal{BT}^{*}_{n}$ with $T^{\prime}$ obtained from $T$ by replacing the pending subtree $T_{v}$ by a subtree $T_{v}^{\prime}$ of the same size. Following [11] again, a binary recursive tree shape statistic of length $x$ is an ordered pair $(\lambda,r)$ consisting of a vector $\lambda\in\mathbb{R}^{2}$ and an $x$-vector of symmetric functions $r_{i}:\mathbb{R}^{2}\times\mathbb{R}^{2}\rightarrow\mathbb{R}$. Here, $x$ is the number of recursions and $\lambda$ with $\lambda_{i}=r_{i}(T)$ for $T$ with $n=1$ contains the start value for each recursion, for $i=1,\ldots,x$. The vector $r$ contains the recursions themselves. Each recursion operates on 2 vectors of length $x$, namely $(r_{1}(T_{a}),\ldots,r_{x}(T_{a}))$ and $(r_{1}(T_{b}),\ldots,r_{2}(T_{b}))$, where $T=(T_{a},T_{b})$ is the standard decomposition of a binary tree $T$.
@@
    where $N_{a}(T)$ and $N_{b}(T)$ are as defined above. Note again that whenever $n=0$, i.e., when we consider $T^{\emptyset}$, the sum of the Colless index is empty, implying we have $C(T^{\emptyset})=0$.
@@
  • the $\Delta_{SC}$ index of a rooted binary tree $T$ is defined as: $\Delta_{CS}(T)=S(T)-C(T)=-\Delta_{CS}$.
@@ 2.2.3 Known results on the Sackin index
The Sackin index is an imbalance index, and it was shown in [10] that $S(T)\geq-2^{k_{n}}+n\cdot(k_{n}+1)$, where $k_{n}=\lceil\log_{2}(n)\rceil$, and that the given bound is tight. The trees that achieve this minimum have been fully characterized in [10], and it was shown there that the set of minimal trees consists precisely of the ones in which each pair of leaves $x$ and $y$ has the property that $|\delta_{x}-\delta_{y}|\leq 1$. In particular, this implies that for all $n\in\mathbb{N}$, $T^{mb}_{n}$ and $T^{gfb}_{n}$ are contained in the set of Sackin-minimal trees. Moreover, from Section 2.2.1 it can be easily seen that if a tree $T=(T_{a},T_{b})$, with $T_{a}\in\mathcal{BT}^{*}_{n_{a}}$ and $T_{b}\in\mathcal{BT}^{*}_{n_{b}}$, is a minimal Sackin tree with $n$ leaves, then ${n}_{a}-{n}_{b}\leq n^{gfb}_{a}-n^{gfb}_{b}$, where $n^{gfb}_{a}$ and $n^{gfb}_{b}$ denote the leaf numbers of the maximal pending subtrees of $T_{n}^{gfb}$, respectively, with $n^{gfb}_{a}\geq n^{gfb}_{b}$. This is due to the fact that otherwise, we could find two leaves $x$, $y$ in $T$ with $|\delta_{x}-\delta_{y}|\geq 2$. Also, for each pair $(n_{a},n_{b})$ with $n_{a}+n_{b}=n$ and $n_{a}-n_{b}\leq n_{a}^{gfb}-n_{b}^{gfb}$ there is a Sackin minimal tree (as these minima can be constructed from attaching cherries to leaves of $T_{k_{n}-1}^{fb}$, cf. [10, 11]). We therefore refer to such pairs $(n_{a},n_{b})$ with $n_{a}+n_{b}=n$ and $n_{a}-n_{b}\leq n_{a}^{gfb}-n_{b}^{gfb}$ as Sackin partitions of $n$.
@@ 2.2.4 Known results on the Colless index
The Colless index is an imbalance index [8, 11]. Several explicit formulas to express $c_{n}$ for each value of $n\in\mathbb{N}$, where $c_{n}$ denotes the minimum value of the Colless index of a tree in $\mathcal{BT}^{*}_{n}$, have been established [13, 8], but we will exemplarily use only the following one: Let $\breve{s}(x)$ denote the so-called triangle wave, i.e., the distance from $x\in\mathbb{R}$ to its nearest integer, which implies $\breve{s}(x)=\min_{z\in\mathbb{Z}}|x-z|$. Then,
@@ 3 Results
It is the main goal of this section to show that the Sackin and Colless indices are merely compound imbalance indices as they are linear combinations of $N_{a}$ and $N_{b}$, which we will show to be (im)balance indices of their own. Along the way, we will also investigate $\Delta_{CS}$ and $\Delta_{SC}$, which will also turn out to be (im)balance indices as they are closely related to $N_{b}$. We start our investigation with $\Delta_{CS}$, $\Delta_{SC}$ and $N_{b}$.
@@ 3.1 $\Delta_{CS}$, $\Delta_{SC}$ and $N_{b}$: three new (im)balance indices on $\mathcal{BT}^{*}_{n}$
Before we can prove this main result, we need some preliminary results. We begin with the following lemma, which provides a simple but important property of this tree shape statistic. In fact, it establishes the close link between $\Delta_{CS}$ and $N_{b}$.
@@
Note that Lemma 2 implies that $\Delta_{CS}$ is maximized whenever $N_{b}$ is minimized and vice versa. Since $\Delta_{CS}=-\Delta_{SC}$, it also shows that $\Delta_{SC}=2N_{b}$, implying that $\Delta_{SC}$ and $N_{b}$ have precisely the same minima and maxima.
Next, we derive a recursion for $\Delta_{CS}$. This will be the crucial ingredient needed in the analysis of the minimum value of $\Delta_{CS}$ as well as the trees that achieve it.
@@
From Lemma 2 we learn that $\Delta_{CS}$ is maximized whenever $N_{b}$ is minimized and vice versa. We now use this fact to derive the maximum value of $\Delta_{CS}$ as well as the trees that achieve it.
@@ 3.1.1 Maximizing $\Delta_{CS}$ and minimizing $N_{b}$ as well as $\Delta_{SC}$
It is the aim of this subsection to show that $T^{cat}_{n}$ is the unique tree maximizing $\Delta_{CS}$ (and minimizing $N_{b}$ as well as $\Delta_{SC}$) and to derive the maximum value of $\Delta_{CS}$ (and minimum value of $N_{b}$ and $\Delta_{SC}$, respectively).
@@
We now turn our attention to the more intricate minimum of $\Delta_{CS}$ and maximum of $\Delta_{SC}$ and $N_{b}$.

### v2

@@ Abstract
(Im)balance indices can be used to quantify the (im)balance of trees by assigning numerical scores to them. An easy way to generate a new index is to construct a compound index, e.g., a linear combination of established indices. Two of the most prominent and widely used imbalance indices are the Sackin index and the Colless index. In this study, we show that these classic indices are themselves compound in nature: they can be decomposed into more elementary components that independently satisfy the defining properties of a tree (im)balance index. We further show that the difference Colless minus Sackin results in another imbalance index that is minimized (amongst others) by all Colless minimal trees. Conversely, the difference Sackin minus Colless forms a balance index. Finally, we compare the building blocks of which the Sackin and the Colless indices consist to these indices as well as to the stairs2 index, which is another index from the literature. Our results suggest that the elementary building blocks we identify are not only foundational to established indices but also valuable tools for analyzing disagreement among indices when comparing the balance of different trees. Along the way, we investigate the so-called echelon tree, which plays an important role for several (im)balance indices, and present the first non-recursive algorithm to construct it.
@@ 1 Introduction
This variability has led to the continuing development of new indices, often with the aim of capturing specific aspects of tree balance more effectively. A common approach in the design of new indices is to combine established ones into compound indices, for instance via simple linear combinations [10, Chapter 25], [16, 3, 22]. Among the most commonly used base indices for such compound construction are the Sackin index [24, 9], which measures the total of all subtree sizes, and the Colless index [6, 7], which measures subtree size imbalances across all internal nodes. Both have been used extensively in theoretical studies and practical applications [1, 2, 28, 10, 15], and they often serve as components in the construction of new indices [10, 3, 16, 22].
@@
Building on this decomposition of the Sackin and Colless indices, we even identify two more indices, which are new to the literature and which turn out to be closely related to $N_{a}$ and $N_{b}$: the difference Colless-Sackin, $\Delta_{CS}$, forms an imbalance index that remains minimal (amongst others) for the same class of balanced trees that minimize Colless, while the difference Sackin-Colless, $\Delta_{SC}$, constitutes a balance index. Concerning all newly discovered indices, i.e., $N_{a}$, $N_{b}$, $\Delta_{CS}$ and $\Delta_{SC}$, we are able to fully characterize their extremal trees for all $n$, and in this characterization, the so-called echelon tree plays a decisive role. The echelon tree has so far mainly been known from the context of two other tree balance indices from the literature, namely the so-called stairs2 and $B_{1}$ balance indices [8, 11]. In our study, we also present some new insight into this tree and present, for instance, the first non-recursive way to generate and define it.
@@ 2.1 Definitions and notation
The most important concept for the present manuscript is that of a rooted binary tree (or simply tree for short), which is a directed graph $T=(V(T),E(T))$, with vertex set $V(T)$ and edge set $E(T)$, containing precisely one vertex of in-degree zero, namely the root (denoted by $\rho$), such that for every $v\in V(T)$ there exists a unique path from $\rho$ to $v$, there are no vertices with out-degree one [^1] Note that this definition of a binary tree (sometimes also called tree-shape or topology) is common in mathematical phylogenetics [25, 10], whereas in computer science, vertices of out-degree one are usually allowed in binary trees [20]. and such that each vertex either has has out-degree 2 (such vertices will be referred to as inner vertices) or 0 (such vertices will be referred to as leaves). For every $n\in\mathbb{N}_{\geq 1}$, we denote by $\mathcal{BT}^{*}_{n}$ the set of (isomorphism classes of) rooted binary trees with $n$ leaves.
@@
Recall that whenever you have a rooted tree $T$, you can define the depth $\delta_{v}$ of a vertex $v$ of $T$ as the length of the unique path from the root $\rho$ of $T$ to $v$. The height $h(T)$ of $T$ is then simply the maximum of all leaf depths, i.e., $h(T)=\max_{x\in V^{1}(T)}\delta_{x}$. Given a subset $X^{\prime}$ of leaves of $T$, the lowest common ancestor or lca of $X^{\prime}$ is the unique ancestor of all leaves in $X^{\prime}$ with the largest depth.
@@
We continue with the greedy from the bottom tree, denoted by $T^{gfb}_{n}$. It is the rooted binary tree with $n$ leaves that results from greedily clustering trees of minimal leaf numbers, starting with $n$ single vertices and proceeding until only one tree is left as described by [7, Algorithm 2].
@@
Another tree that is of high relevance in our manuscript is the binary echelon tree or echelon tree for short, which is denoted by $T_{n}^{be}$. This tree was recursively defined in [8] as follows: If $n=0$, $T_{n}^{be}$ is the empty tree $T^{\emptyset}$, and if $n=1$, $T_{n}^{be}$ consists of a single vertex. Now, for $n>1$, let $k_{n}=\left\lceil\log_{2}n\right\rceil$. Then, $T_{n}^{be}$ contains a maximal pending subtree $T_{a}$ which equals $T_{k_{n}-1}^{fb}$ and a maximal pending subtree $T_{b}$ which equals $T_{n-2^{k_{n}-1}}^{be}$. Thus, the binary echelon tree is uniquely defined (up to isomorphism) for all $n$. See Figure 4 for an exemplary depiction of $T_{9}^{be}$ and a comparison to $T_{9}^{gfb}$.
@@
In order to define (im)balance indices, however, we first need to consider tree shape statistics: A tree shape statistic is a function $t:\mathcal{BT}^{*}_{n}\rightarrow\mathbb{R}$. Based on this, a tree imbalance index is defined as follows: A tree shape statistic $t$ is called an imbalance index if and only if
@@
Similarly, a tree shape statistic $t$ is called a balance index if and only if
@@
Following, for instance, [10], we call a balance or imbalance index local if it fulfills the following criterion: If two trees differ only in one rooted subtree, the difference between their indices is equal to the difference between the indices of these differing subtrees [10, 23]. More precisely, a balance or imbalance index $t$ is local, if $t(T)-t(T^{\prime})=t(T_{v})-t(T_{v}^{\prime})$ for $T$, $T^{\prime}\in\mathcal{BT}^{*}_{n}$ with $T^{\prime}$ obtained from $T$ by replacing the pending subtree $T_{v}$ by a subtree $T_{v}^{\prime}$ of the same size. Following [10, 21], a binary recursive tree shape statistic of length $x$ is an ordered pair $(\lambda,r)$ consisting of a vector $\lambda\in\mathbb{R}^{x}$ and an $x$-vector $r$ of symmetric functions $r_{i}:\mathbb{R}^{x}\times\mathbb{R}^{x}\rightarrow\mathbb{R}$, also called recursions. Here, $x$ is called length of $r$ and denotes the number of recursions. Moreover, $\lambda$ with $\lambda_{i}=r_{i}(T)$ for $T$ with $n=1$ contains the base case for each recursion, for $i=1,\ldots,x$. Note that even though each recursion $r_{i}$ operates on 2 vectors of length $x$, namely $(r_{1}(T_{a}),\ldots,r_{x}(T_{a}))$ and $(r_{1}(T_{b}),\ldots,r_{x}(T_{b}))$, that contain real numbers, following [10] we use the shorthand $r_{i}(T_{a},T_{b})$ instead of $r_{i}((r_{1}(T_{a}),\ldots,r_{x}(T_{a})),(r_{1}(T_{b}),\ldots,r_{x}(T_{b})))$. Here, $T=(T_{a},T_{b})$ is the standard decomposition of a binary tree $T\in\mathcal{BT}^{*}_{n}$.
@@
    where $N_{a}(T)$ and $N_{b}(T)$ are as defined above. Note again that whenever $n=0$, i.e., when we consider $T^{\emptyset}$, the sum of the Colless index is empty, implying that we have $C(T^{\emptyset})=0$.
@@
  • the $\Delta_{SC}$ index of a rooted binary tree $T$ is defined as: $\Delta_{SC}(T)=S(T)-C(T)=-\Delta_{CS}(T)$.
@@ 2.2.3 Known results on the Sackin index
The Sackin index is an imbalance index, and it was shown in [9] that $S(T)\geq-2^{k_{n}}+n\cdot(k_{n}+1)$, where $k_{n}=\lceil\log_{2}(n)\rceil$, and that the given bound is tight. The trees that achieve this minimum have been fully characterized in [9], and it was shown there that the set of minimal trees consists precisely of the ones in which each pair of leaves $x$ and $y$ has the property that $|\delta_{x}-\delta_{y}|\leq 1$. In particular, this implies that for all $n\in\mathbb{N}$, $T^{mb}_{n}$ and $T^{gfb}_{n}$ are contained in the set of Sackin-minimal trees. Moreover, from Section 2.2.1 it can be easily seen that if a tree $T=(T_{a},T_{b})$, with $T_{a}\in\mathcal{BT}^{*}_{n_{a}}$ and $T_{b}\in\mathcal{BT}^{*}_{n_{b}}$, is a minimal Sackin tree with $n$ leaves, then ${n}_{a}-{n}_{b}\leq n^{gfb}_{a}-n^{gfb}_{b}$, where $n^{gfb}_{a}$ and $n^{gfb}_{b}$ denote the leaf numbers of the maximal pending subtrees of $T_{n}^{gfb}$, respectively, with $n^{gfb}_{a}\geq n^{gfb}_{b}$. This is due to the fact that otherwise, we could find two leaves $x$, $y$ in $T$ with $|\delta_{x}-\delta_{y}|\geq 2$. Also, for each pair $(n_{a},n_{b})$ with $n_{a}+n_{b}=n$ and $n_{a}-n_{b}\leq n_{a}^{gfb}-n_{b}^{gfb}$ there is a Sackin minimal tree $T$ with $T=(T_{a},T_{b})$ and $T_{a}$ of size $n_{a}$ and $T_{b}$ of size $n_{b}$ (as these minima can be constructed from attaching cherries to leaves of $T_{k_{n}-1}^{fb}$, cf. [9, 10]). We therefore refer to such pairs $(n_{a},n_{b})$ with $n_{a}+n_{b}=n$ and $n_{a}-n_{b}\leq n_{a}^{gfb}-n_{b}^{gfb}$ as Sackin partitions of $n$.
@@ 2.2.4 Known results on the Colless index
The Colless index is an imbalance index [7, 10]. Several explicit formulas to express $c_{n}$ for each value of $n\in\mathbb{N}$, where $c_{n}$ denotes the minimum value of the Colless index of a tree in $\mathcal{BT}^{*}_{n}$, have been established [13, 7], but we will exemplarily use only the following one: Let $\breve{s}(x)$ denote the distance from $x\in\mathbb{R}$ to its nearest integer, which implies $\breve{s}(x)=\min_{z\in\mathbb{Z}}|x-z|$. Then,
@@ 3 Results
The main goal of this section is twofold: First, we want to show that the Sackin and Colless indices are merely compound imbalance indices as they are linear combinations of $N_{a}$ and $N_{b}$, which we will show to be (im)balance indices of their own. Along the way, we will also investigate $\Delta_{CS}$ and $\Delta_{SC}$, which will also turn out to be (im)balance indices as they are closely related to $N_{b}$. Second, we will show that the echelon tree plays an important role for three of these new (im)balance indices, and we will investigate its properties further in Section 3.3. We start our investigation with $\Delta_{CS}$, $\Delta_{SC}$ and $N_{b}$.
@@ 3.1 $\Delta_{CS}$, $\Delta_{SC}$ and $N_{b}$: three new (im)balance indices on $\mathcal{BT}^{*}_{n}$
Before we can prove this main result, we need some preliminary results. We begin with the following lemma, which provides a simple but important property of $\Delta_{CS}$. In fact, it establishes the close link between $\Delta_{CS}$ and $N_{b}$.
@@
Note that Lemma 2 implies that $\Delta_{CS}$ is maximized whenever $N_{b}$ is minimized and vice versa. Since $\Delta_{CS}=-\Delta_{SC}$, it also shows that $\Delta_{SC}=2N_{b}$, implying that $\Delta_{SC}$ and $N_{b}$ have precisely the same minima and maxima. Therefore, in the following we mainly focus on $\Delta_{CS}$. We only directly mention $\Delta_{SC}$ and $N_{b}$ in the main theorems or whenever we give specific values.
Next, we derive a recursion for $\Delta_{CS}$. This will be one of the most important ingredients needed in the analysis of the minimum value of $\Delta_{CS}$ as well as the trees that achieve it.
@@ 3.1.1 Maximizing $\Delta_{CS}$ and minimizing $N_{b}$ as well as $\Delta_{SC}$
It is the aim of this subsection to show that $T^{cat}_{n}$ is the unique tree maximizing $\Delta_{CS}$ and to derive the maximum value of $\Delta_{CS}$.
@@
We now turn our attention to the more intricate minimum of $\Delta_{CS}$.